# Fluid Mechanics Unit Italian Aerospace Research Center (CIRA) Capua (CE), Italy

### v1

@@ 4 A class of finite-difference entropy-conserving methods
$=$ $\int_{\Omega}\left[\left[c_{v}\cdot\log\frac{e}{e_{r}}-c_{v}(\gamma-1)\cdot\log\frac{\rho}{\rho_{r}}-\gamma c_{v}\right]\cdot\bm{\mathcal{C}}_{\rho}+\frac{c_{v}}{e}\cdot\left(\bm{\mathcal{C}}_{\rho e}+\bm{\mathcal{P}}_{\rho e}\right)\right]$
$=$ $\underbrace{c_{v}\int_{\Omega}\left[\frac{1}{e}\cdot\bm{\mathcal{P}}_{\rho e}-(\gamma-1)\log\frac{\rho}{\rho_{r}}\cdot\bm{\mathcal{C}}_{\rho}\right]}_{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(1)}+\underbrace{c_{v}\int_{\Omega}\left[\left(\log\frac{e}{e_{r}}-\gamma\right)\cdot\bm{\mathcal{C}}_{\rho}+\frac{1}{e}\cdot\bm{\mathcal{C}}_{\rho e}\right]}_{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}(2)}$
In terms of variable units, the explicit normalization of internal energy and density by their reference quantities within the logarithmic expressions of entropy have now been included for clarity of presentation. As an example, we show this below in Eqs. (54) and (55) for the one-dimensional case assuming a set of biased operators (note that the same mechanics apply for the opposite biasing and for the central stencils as well). The integral notation used below actually implies summation across nodal points in the discrete setting according to the SBP diagonal norm that encodes the interface flux spacings [41]
$\underbrace{c_{v}\int_{\Omega}\left[\frac{1}{e}\cdot\bm{\mathcal{P}}^{(+)}_{\rho e}-(\gamma-1)\log\frac{\rho}{\rho_{r}}\cdot\bm{\mathcal{C}}^{(+)}_{\rho}\right]}_{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(1)}:$ $c_{v}\int_{\Omega}\left[e^{-1}p\cdot\delta_{x}^{+}u-(\gamma-1)\log\frac{\rho}{\rho_{r}}\cdot\delta_{x}^{+}(\tilde{\rho}u)\right]$ (54)
$=c_{v}(\gamma-1)\int_{\Omega}\left[\rho\cdot\delta_{x}^{+}u-\log\frac{\rho}{\rho_{r}}\cdot\delta_{x}^{+}(\tilde{\rho}u)\right]$
$=c_{v}(\gamma-1)\left[\rho u-\log\frac{\rho}{\rho_{r}}\cdot\tilde{\rho}u\right]\big{|}_{\partial\Omega}$
$-c_{v}(\gamma-1)\int_{\Omega}u\underbrace{\left[\delta_{x}^{-}\rho-\tilde{\rho}\cdot\delta_{x}^{-}\left(\log\frac{\rho}{\rho_{r}}\right)\right]}_{\text{need}\ =0}$
$\underbrace{c_{v}\int_{\Omega}\left[\left(\log\frac{e}{e_{r}}-\gamma\right)\cdot\bm{\mathcal{C}}^{(+)}_{\rho}+\frac{1}{e}\cdot\bm{\mathcal{C}}^{(+)}_{\rho e}\right]}_{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}(2)}:$ $c_{v}\int_{\Omega}\left[\left(\log\frac{e}{e_{r}}-\gamma\right)\cdot\delta_{x}^{+}(\tilde{\rho}u)+e^{-1}\cdot\delta_{x}^{+}(\tilde{\rho}u\tilde{e})\right]$ (55)
$=c_{v}\left[\left(\log\frac{e}{e_{r}}-\gamma\right)\cdot(\tilde{\rho}u)+e^{-1}\cdot(\tilde{\rho}u\tilde{e})\right]\big{|}_{\partial\Omega}$
$-c_{v}\int_{\Omega}\tilde{\rho}u\underbrace{\left[-\delta_{x}^{-}\left(\log\frac{e^{-1}}{e_{r}^{-1}}\right)+\tilde{e}\cdot\delta_{x}^{-}e^{-1}\right]}_{\text{need}\ =0}$
@@
$\bm{\mathcal{C}}_{\rho s}\big{|}_{\ell}\equiv\mathcal{F}_{\rho s,\ell}$ $=$ $\overbrace{(\tilde{\rho}u)_{\ell}\cdot\left(c_{v}\left(\log\frac{e}{e_{r}}\right)_{\ell}-c_{v}(\gamma-1)\left(\log\frac{\rho}{\rho_{r}}\right)_{\ell}\right)}^{(\rho us)_{\ell}}$
$+\ \underbrace{\gamma c_{v}\left(\rho_{\ell}u_{\ell}-(\tilde{\rho}u)_{\ell}\right)-c_{v}\left(\rho_{\ell}u_{\ell}-(\tilde{\rho}u\tilde{e})_{\ell}(e^{-1})_{\ell}\right)}_{O(\Delta x^{p})}$
Equation (4) is valid for all of the SBP stencils—whether central of forward/backward biased. We further note that in cases where $\tilde{\rho}\neq\rho$ and $\tilde{e}\neq e$ (as will be the case for the resulting EC schemes) then the resulting boundary flux includes a persistent zero-consistent component.
The spurious entropy volumetric terms that result from the summation-by-parts manipulations in Eqs. (54) and (55) are canceled point-wise by selecting the auxiliary variables $\tilde{\rho}\equiv\frac{\delta_{x}^{-}\rho}{\delta_{x}^{-}\log\rho}\triangleq\overline{\rho}^{(\log,-)}$ and $\tilde{e}\equiv\frac{\delta_{x}^{-}\log e^{-1}}{\delta_{x}^{-}e^{-1}}\triangleq\overline{e}^{(H\log,-)}$. Per the discussion of Section 3, these new variables constitute specialized averages of the density and internal energy. The resulting method is then successfully entropy conserving and can also be associated with a local entropy-conserving Tadmor-type flux condition (see Remark 2). In addition to being entropy conserving, it is also kinetic-energy-preserving by construction and can be shown to be pressure-equilibrium-preserving (PEP) as discussed in Section 4.1. The finite difference representation that is induced in the discrete entropy equation can be identified from the structure of the boundary flux in Eq. (4). Specifically, the forward biased rendition gives
@@
$\ +\underbrace{\left[c_{v}(\gamma-1)\cdot\left(\rho\delta^{+}_{x}u+u\delta^{-}_{x}\rho\right)-\gamma c_{v}\cdot\delta_{x}^{+}(\tilde{\rho}u)+c_{v}\left(e^{-1}\delta_{x}^{+}(\tilde{\rho}u\tilde{e})+\tilde{\rho}u\tilde{e}\delta_{x}^{-}e^{-1}\right)\right]}_{O(\Delta x)^{p}}\ .$
Thus we recover a split form in the entropy dynamics that is based on the special nonlinear averages $\tilde{\rho}$ and $\tilde{e}$; and besides the main entropy terms, there is furthermore a zero-consistent correction. Note that analogous expressions arise when switching the biasing and also when using the central operators (not shown).
@@
The point-wise cancellations of the spurious dynamics in the discrete entropy equation (see Eqs.→(54)–(55)) suggest that the FD-EC schemes also satisfy the following Tadmor-type EC flux condition: $\sum_{m=1}^{2+d}(w_{m,i+1}-w_{m,i})\mathcal{F}^{(n)}_{m,i+1/2}=(\tilde{\psi}^{(n)}_{i+1}-\tilde{\psi}^{(n)}_{i})$ for $i=1,\dots,N-1$, where $\tilde{\psi}^{(n)}_{1}\equiv\tilde{\psi}^{(n)}_{\ell}=(\sum_{m}^{2+d}w_{m,1}\mathcal{F}^{(n)}_{m,\ell}-\mathcal{F}^{(n)}_{\rho s,\ell})$ (see Section 3.2.1 in [41]).
@@
The specialized auxiliary variables $\tilde{\rho}$ and $\tilde{e}$ need to be calculated with respect to each directional derivative. For example, fluxes in the $n$ direction would require calculating $\widetilde{\rho}^{(\pm,n)}\triangleq\frac{\delta_{x_{n}}^{\pm}\rho}{\delta_{x_{n}}^{\pm}\log\rho}\equiv\overline{\rho}^{(\log,\pm,n)}$ and $\widetilde{e}^{(\pm,n)}\triangleq\frac{\delta_{x_{n}}^{\pm}\log e^{-1}}{\delta_{x_{n}}^{\pm}e^{-1}}\equiv\overline{e}^{(H\log,\pm,n)}$ in the biased case.
@@
$\mathcal{F}_{\ell}^{(n)}=\left[\begin{array}[]{c}\mathcal{F}_{\rho}\\ \\ \mathcal{F}_{\rho u_{m}}\\ \\ \mathcal{F}_{\rho E}\end{array}\right]_{\ell}^{(n)}=\left[\begin{array}[]{c}(\tilde{\rho}u_{n})_{\ell}\\ \\ \frac{1}{2}(\tilde{\rho}u_{n}u_{m})_{\ell}+\frac{1}{2}(\tilde{\rho}u_{n})_{\ell}(u_{m})_{\ell}+\delta_{mn}\cdot p_{\ell}\\ \\ \frac{1}{2}\sum_{m=1}^{d}(u_{m})_{\ell}(\tilde{\rho}u_{n}u_{m})_{\ell}+(\tilde{\rho}u_{n}\tilde{e})_{\ell}+(u_{n})_{\ell}p_{\ell}\end{array}\right]$ (68)
@@
It is insightful to also observe the structure of the resulting fluxes on the domain interior. For example, the two-point central difference stencils give the following “wide-width" interior interface fluxes
@@
and the two-point biased difference stencils yield the following parameterized “narrow-width" interior interface fluxes
$\alpha\cdot\left(\bm{\mathcal{C}}^{(n,+)}+\bm{\mathcal{P}}^{(n,+)}\right)_{\text{2pt}}+(1-\alpha)\cdot\left(\bm{\mathcal{C}}^{(n,-)}+\bm{\mathcal{P}}^{(n,-)}\right)_{\text{2pt}}$
$\to\ \mathcal{F}^{(n)}_{i+1/2}\big{|}^{\alpha}_{\text{narrow}}=\left[\begin{array}[]{c}\overline{(\rho_{i},\rho_{i+1})}^{\log,n}\left((1-\alpha)\cdot u_{n,i}+\alpha\cdot u_{n,i+1}\right)\\ \\ \mathcal{F}_{\rho,i+1/2}^{(n)}\cdot\frac{(u_{m,i+1}+u_{m,i})}{2}+\delta_{mn}\cdot\left((1-\alpha)\cdot p_{i+1}+\alpha\cdot p_{i}\right)\\ \\ \mathcal{F}_{\rho,i+1/2}^{(n)}\cdot\left(\sum_{m=1}^{d}\frac{u_{m,i}u_{m,i+1}}{2}+\overline{(e_{i},e_{i+1})}^{H\log,n}\right)\\ +\ (1-\alpha)\cdot u_{n,i}p_{i+1}+\alpha\cdot u_{n,i+1}p_{i}\end{array}\right]$ (84)
@@
Some interesting insights transpire from studying the interior flux forms stemming from the two-point schemes. First, we note that the central stencils induce a wider node dependence and furthermore that their fluxes couple the thermodynamic interface reconstructions with that of the transport velocity. On the other hand, the biased stencils—at least in the two-point scenario—split up these variable reconstructions. Next, we note that the biased stencils yield counter-balanced biasing in the flux with respect to velocity and pressure when $\alpha\neq 1/2$. The associated interface reconstructions in such cases is of lower order than the symmetric case ($\alpha=1/2$, as expected by the fact that such biased stencils are typically one order lower than their centered counterparts. To the authors’ knowledge, this is the first time that EC fluxes with biased interior quantities have been presented, and leveraging such properties toward improving the scheme resolution of sharp gradients constitutes a line of future research.
Also notable upon studying these interior fluxes is the fact that a symmetric combination of the biased two-point operators with $\alpha=1/2$ recovers the second-order flux of Ranocha [25, 26], therefore formalizing its finite-difference compatibility. This exact congruence between the symmetric two-point narrow-width schemes and the Ranocha flux is not expected to hold to high-order, however. Namely, the Ranocha flux (as well as other two-point EC fluxes in the literature [51, 52]) is typically extended to high-order via flux-differencing [16, 41], which is a Richardson-type extrapolation of the low-order two-point fluxes. This, however, need not be the case for the current FD-EC methods which can technically be extended to high-order simply by considering multi-point FD stencils in the definitions of $\tilde{\rho}$ and $\tilde{e}$ as well as the flux derivatives. This is exemplified most simply when inspecting the mass flux term, which motivates the general uniqueness of the respective representations at high order. Assuming a $K$-width central stencil with $p$-point biasing for the associated dual pair schemes, one would have
@@
(via current FD-EC symmetric narrow, $\alpha=1/2$):
@@
[…]
Note that Eqs. (4) and (4) are still expressed in a flux-difference flavor (albeit with multi-point information), which stems naturally from their finite difference origins. Due to the nonlinearity of the logarithmic mean, it is clear that the standard flux-differenced implementation of Ranocha’s method will differ from the newly-presented symmetric narrow-width method for $K\neq 1$, as the latter would utilize a multi-point logarithmic means of density rather than two-point logarithmic means. Suitable algorithms for handling such multi-point nonlinear averages near singular points (i.e., where the quotient formulas from Section 2 are near zero) are still under development; therefore, the numerical results in the following section will focus on the class of two-point FD-EC methods. Appendix A, however, provides an alternate error-reducing treatment for the narrow biased schemes based on flux differencing with multi-point difference stencils.

### v2

@@ 4 A class of finite-difference entropy-conserving methods
$=$ $\int_{\Omega}\left[\left(c_{v}\cdot\log\frac{e}{e_{\text{ref}}}-c_{v}(\gamma-1)\cdot\log\frac{\rho}{\rho_{\text{ref}}}-\gamma c_{v}\right)\cdot\bm{\mathcal{C}}_{\rho}+\frac{c_{v}}{e}\cdot\left(\bm{\mathcal{C}}_{\rho e}+\bm{\mathcal{P}}_{\rho e}\right)\right]$
$=$ $\underbrace{c_{v}\int_{\Omega}\left[\frac{1}{e}\cdot\bm{\mathcal{P}}_{\rho e}-(\gamma-1)\log\frac{\rho}{\rho_{\text{ref}}}\cdot\bm{\mathcal{C}}_{\rho}\right]}_{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(1)}+\underbrace{c_{v}\int_{\Omega}\left[\left(\log\frac{e}{e_{\text{ref}}}-\gamma\right)\cdot\bm{\mathcal{C}}_{\rho}+\frac{1}{e}\cdot\bm{\mathcal{C}}_{\rho e}\right]}_{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}(2)}$
The integral notation used actually implies summation across nodal points in the discrete setting according to the SBP diagonal norm that encodes the interface flux spacings [52]. In terms of variable units, the explicit normalization of internal energy and density by their reference quantities within the logarithmic expressions of entropy have now been included for clarity of presentation. As an example, we show this below in Eqs. (54) and (55) for the one-dimensional case assuming a set of biased operators (note that the same mechanics apply for the opposite biasing and for the central stencils as well). Appendix A shows the analogous entropy-conservation proofs with SBP operators and matrix-vector notation.
$\underbrace{c_{v}\int_{\Omega}\left[\frac{1}{e}\cdot\bm{\mathcal{P}}^{(+)}_{\rho e}-(\gamma-1)\log\frac{\rho}{\rho_{\text{ref}}}\cdot\bm{\mathcal{C}}^{(+)}_{\rho}\right]}_{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(1)}:$ $c_{v}\int_{\Omega}\left[e^{-1}p\cdot\delta_{x}^{+}u-(\gamma-1)\log\frac{\rho}{\rho_{\text{ref}}}\cdot\delta_{x}^{+}(\tilde{\rho}u)\right]$ (54)
$=c_{v}(\gamma-1)\int_{\Omega}\left[\rho\cdot\delta_{x}^{+}u-\log\frac{\rho}{\rho_{\text{ref}}}\cdot\delta_{x}^{+}(\tilde{\rho}u)\right]$
$=c_{v}(\gamma-1)\left[\rho u-\log\frac{\rho}{\rho_{\text{ref}}}\cdot\tilde{\rho}u\right]\big|_{\partial\Omega}$
$-c_{v}(\gamma-1)\int_{\Omega}u\underbrace{\left[\delta_{x}^{-}\rho-\tilde{\rho}\cdot\delta_{x}^{-}\left(\log\frac{\rho}{\rho_{\text{ref}}}\right)\right]}_{\text{need}\ =0}$
[…]
$=c_{v}\left[\left(\log\frac{e}{e_{\text{ref}}}-\gamma\right)\cdot(\tilde{\rho}u)+e^{-1}\cdot(\tilde{\rho}u\tilde{e})\right]\big|_{\partial\Omega}$
$-c_{v}\int_{\Omega}\tilde{\rho}u\underbrace{\left[-\delta_{x}^{-}\left(\log\frac{e^{-1}}{e_{\text{ref}}^{-1}}\right)+\tilde{e}\cdot\delta_{x}^{-}e^{-1}\right]}_{\text{need}\ =0}$
@@
$\bm{\mathcal{C}}_{\rho s}\big|_{\ell}\equiv\mathcal{F}_{\rho s,\ell}=\overbrace{(\tilde{\rho}u)_{\ell}\cdot\left(c_{v}\left(\log\frac{e}{e_{\text{ref}}}\right)_{\ell}-c_{v}(\gamma-1)\left(\log\frac{\rho}{\rho_{\text{ref}}}\right)_{\ell}\right)}^{(\tilde{\rho}u)_{\ell}s_{\ell}\ \sim\ (\rho us)_{\ell}}+\underbrace{\gamma c_{v}\left(\rho_{\ell}u_{\ell}-(\tilde{\rho}u)_{\ell}\right)-c_{v}\left(\rho_{\ell}u_{\ell}-(\tilde{\rho}u\tilde{e})_{\ell}(e^{-1})_{\ell}\right)}_{O(\Delta x^{p})}$
Equation (4) is valid for all of the SBP stencils—whether central or forward/backward biased. We further note that in cases where $\tilde{\rho}\neq\rho$ and $\tilde{e}\neq e$ (as will be the case for the resulting EC schemes) then the resulting boundary flux includes a persistent zero-consistent component.
The spurious entropy volumetric terms that result from the summation-by-parts manipulations in Eqs. (54) and (55) are cancelled point-wise by selecting the auxiliary variables $\tilde{\rho}\equiv\frac{\delta_{x}^{-}\rho}{\delta_{x}^{-}\log\rho}\triangleq\overline{\rho}^{(\log,-)}$ and $\tilde{e}\equiv\frac{\delta_{x}^{-}\log e^{-1}}{\delta_{x}^{-}e^{-1}}\triangleq\overline{e}^{(H\log,-)}$. Per the discussion of Section 3, these new variables constitute specialized averages of the density and internal energy. The resulting method is then successfully entropy conserving and can also be associated with a local entropy-conserving Tadmor-type flux condition (see Remark 2). The finite-difference representation that is induced in the discrete entropy equation can be identified from the structure of the boundary flux in Eq. (4). Specifically, the forward biased rendition gives
@@
$\ +\underbrace{\left[c_{v}(\gamma-1)\cdot\left(\rho\delta^{+}_{x}u+u\delta^{-}_{x}\rho\right)-\gamma c_{v}\cdot\delta_{x}^{+}(\tilde{\rho}u)+c_{v}\left(e^{-1}\delta_{x}^{+}(\tilde{\rho}u\tilde{e})+\tilde{\rho}u\tilde{e}\delta_{x}^{-}e^{-1}\right)\right]}_{O(\Delta x^{p})}\ .$
Thus we recover a split form in the entropy dynamics that is based on the special nonlinear averages $\tilde{\rho}$ and $\tilde{e}$ Besides the main entropy terms, there are also zero-consistent and telescoping (i.e., conservative) corrective contributions. Note that analogous expressions arise when switching the biasing and also when using the central operators (not shown).
@@
The point-wise cancellations of the spurious dynamics in the discrete entropy equation (see Eqs. (54)–(55)) suggest that the FD-EC schemes also satisfy the following Tadmor-type EC flux condition: $\sum_{m=1}^{2+d}(w_{m,i+1}-w_{m,i})\mathcal{F}^{(n)}_{m,i+1/2}=(\tilde{\psi}^{(n)}_{i+1}-\tilde{\psi}^{(n)}_{i})$ for $i=1,\dots,N-1$, where $\tilde{\psi}^{(n)}_{1}\equiv\tilde{\psi}^{(n)}_{\ell}=(\sum_{m}^{2+d}w_{m,1}\mathcal{F}^{(n)}_{m,\ell}-\mathcal{F}^{(n)}_{\rho s,\ell})$ (see Section 3.2.1 in [52]).
@@
The specialized auxiliary variables $\tilde{\rho}$ and $\tilde{e}$ need to be calculated with respect to each directional derivative. For example, fluxes in the $n$ direction would require calculating $\widetilde{\rho}\triangleq\frac{\delta_{x_{n}}^{\pm}\rho}{\delta_{x_{n}}^{\pm}\log\rho}\equiv\overline{\rho}^{(\log,n,\pm)}$ and $\widetilde{e}\triangleq\frac{\delta_{x_{n}}^{\pm}\log e^{-1}}{\delta_{x_{n}}^{\pm}e^{-1}}\equiv\overline{e}^{(H\log,n,\pm)}$ in the biased case.
@@
$\mathcal{F}_{\ell}^{(n)}=\left[\begin{array}[]{c}\mathcal{F}_{\rho}\\ \\ \mathcal{F}_{\rho u_{m}}\\ \\ \mathcal{F}_{\rho E}\end{array}\right]_{\ell}^{(n)}=\left[\begin{array}[]{c}(\tilde{\rho}u_{n})_{\ell}\\ \\ \frac{1}{2}(\tilde{\rho}u_{n}u_{m})_{\ell}+\frac{1}{2}(\tilde{\rho}u_{n})_{\ell}(u_{m})_{\ell}+\delta_{mn}\cdot p_{\ell}\\ \\ \frac{1}{2}\sum_{m=1}^{d}(u_{m})_{\ell}(\tilde{\rho}u_{n}u_{m})_{\ell}+(\tilde{\rho}u_{n}\tilde{e})_{\ell}+(u_{n})_{\ell}p_{\ell}\end{array}\right].$ (68)
@@
It is insightful to also observe the structure of the resulting fluxes on the domain interior. For example, the two-point central-difference stencils give the following “wide-width" interior interface fluxes
@@
and the two-point biased difference stencils yield the following parameterized “narrow-width" interior interface fluxes:
$\omega\cdot\left(\bm{\mathcal{C}}^{(n,+)}+\bm{\mathcal{P}}^{(n,+)}\right)_{\text{2pt}}+(1-\omega)\cdot\left(\bm{\mathcal{C}}^{(n,-)}+\bm{\mathcal{P}}^{(n,-)}\right)_{\text{2pt}}$
$\to\ \mathcal{F}^{(n)}_{i+1/2}\big|^{\omega}_{\text{narrow}}=\left[\begin{array}[]{c}\overline{(\rho_{i},\rho_{i+1})}^{\log,n}\left((1-\omega)\cdot u_{n,i}+\omega\cdot u_{n,i+1}\right)\\ \\ \mathcal{F}_{\rho,i+1/2}^{(n)}\cdot\frac{(u_{m,i+1}+u_{m,i})}{2}+\delta_{mn}\cdot\left((1-\omega)\cdot p_{i+1}+\omega\cdot p_{i}\right)\\ \\ \mathcal{F}_{\rho,i+1/2}^{(n)}\cdot\left(\sum_{m=1}^{d}\frac{u_{m,i}u_{m,i+1}}{2}+\overline{(e_{i},e_{i+1})}^{H\log,n}\right)\\ +\ (1-\omega)\cdot u_{n,i}p_{i+1}+\omega\cdot u_{n,i+1}p_{i}\end{array}\right].$ (84)
@@
Some interesting insights transpire from studying the interior flux forms stemming from the two-point schemes. First, we note that the central stencils induce a wider node dependence and furthermore that their fluxes couple the thermodynamic interface reconstructions with that of the transport velocity. On the other hand, the biased stencils—at least in the two-point scenario—split up these variable reconstructions. Next, we note that the biased stencils yield counter-balanced biasing in the flux with respect to velocity and pressure when $\omega\neq 1/2$. The associated interface reconstructions in such cases are of lower order than the symmetric case ($\omega=1/2$), as expected by the fact that such biased stencils are typically one order lower than their centered counterparts. To the authors’ knowledge, this is the first time that EC fluxes with biased interior quantities have been presented, and leveraging such properties toward improving the scheme resolution of sharp gradients constitutes a line of future research.
Also notable upon studying these interior fluxes is the fact that a symmetric combination of the biased two-point operators with $\omega=1/2$ recovers the second-order flux of Ranocha [30, 31], therefore formalizing its finite-difference compatibility. This exact congruence between the symmetric two-point narrow-width schemes and the Ranocha flux is not expected to hold to high-order, however. Namely, the Ranocha flux (as well as other two-point EC fluxes in the literature [15, 62]) is typically extended to high-order via flux-differencing [22, 52], which is a Richardson-type extrapolation of the low-order two-point fluxes. This, however, need not be the case for the current FD-EC methods which can technically be extended to high-order simply by considering multi-point FD stencils in the definitions of $\tilde{\rho}$ and $\tilde{e}$ as well as the flux derivatives. This is exemplified most simply when inspecting the mass flux term, which motivates the general uniqueness of the respective representations at high order. Assuming a $K$-width central stencil with $p$-point biasing for the associated dual pair schemes, one would have
@@
(via current FD-EC symmetric narrow, $\omega=1/2$):
@@
$\text{(on the interior)}:\quad d_{i,i-k}^{\text{c}}=-d_{i,i+k}^{\text{c}}\ \text{and}\ d_{i,i}^{\text{c}}=0\ .$ (89)
Note that Eqs. (4) and (4) are still expressed in a flux-difference flavor (albeit with multi-point information), which stems naturally from their finite difference origins. Due to the nonlinearity of the logarithmic mean, it is clear that the standard flux-differenced implementation of Ranocha’s method will differ from the newly presented symmetric narrow-width method for $K\neq 1$, as the latter would utilize a multi-point logarithmic filters of density rather than two-point logarithmic means. Suitable algorithms are still under development for handling such multi-point nonlinear averages near singular points (i.e., where the quotient formulas from Section 2 are near zero) and to properly enforce positivity of the output; therefore, the numerical results in the following section will focus on the class of two-point FD-EC methods. Appendix D, however, provides an alternate error-reducing treatment for the narrow biased schemes based on flux differencing with multi-point difference stencils.